# Contributions of $\rho(770,1450)\to\omega\pi$ for the Cabibbo-favored $D\to h\omega\pi$ decays

### v1

@@ III Numerical results and Discussions
The crucial input in this work is the coupling constant $g_{\rho\omega\pi}$ in Eq. (21) for the subprocess $\rho(770)\to\omega\pi$; its value can be estimated with the relation $g_{\rho\omega\pi}\approx 3g_{\rho\pi\pi}^{2}/(8\pi^{2}F_{\pi})$ prc83.048201 with $F_{\pi}=92$ MeV PDG-2024 and can also be calculated from the decay width of $\omega\to\pi^{0}\gamma$ PDG-2024 . In the numerical calculation of this work, we adopt the value $g_{\rho\omega\pi}=16.0\pm 2.0$ GeV-1 JHEP2401-047 by taking into account the corresponding fitted and theoretical values in Refs. prd94.112001 ; prd86.057301 ; prd61.072003 ; JETPL94-734 ; prd55.249 ; prd77.113011 ; prd86.037302 ; prd88.054013 for it. In view of the expression of $F_{\omega\pi}(s)$, we have a constraint JHEP2401-047
@@
Table 3: The predicted branching fractions for the relevant quasi-two-body $D_{(s)}$ decays with the subprocesses $\rho(770,1450)^{+}\to\omega\pi^{+}$.
@@
The absolute branching fraction for the three-body decay $D^{+}\to K^{0}_{S}\pi^{+}\omega$ was determined to be $(0.707\pm 0.041_{\rm stat}\pm 0.029_{\rm syst})\%$ by BESIII in Ref. prd105.032009 , which is $4.6$ times larger than the prediction $(1.53\pm 0.37)\times 10^{-3}$ in Table 3 for the quasi-two-body decay $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$. We need to stress that the contribution weight of $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ in the corresponding three-body decay could be enhanced by improving the calculation method employed in this work, since our prediction for $D^{+}\to K^{0}_{S}[\rho(770)^{+}\to]\pi^{+}\pi^{0}$ in Table 2 is just around half of the data in PDG-2024 . Which also means the branching fraction for $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ could be twice of the corresponding prediction in Table 3. In addition, the $D^{+}\to b_{1}(1235)$ transition with a emitted $K^{0}_{S}$, could generate the $\omega\pi^{+}$ pair from $b_{1}(1235)$ for $D^{+}\to K^{0}_{S}\pi^{+}\omega$ decay. The $\omega\pi^{+}$ pair generated from the $D^{+}\to b_{1}(1235)$ transition is not like it from the matrix element $\langle\omega\pi|V^{\mu}\!-\!A^{\mu}|0\rangle$ in $D_{s}^{+}\to\eta\omega\pi^{+}$ decay, whose amplitude is just proportional to the mass difference between $u$ and $d$ quarks. The measurement of ${\mathcal{B}}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e})\times{\mathcal{B}}(b_{1}(1235)^{0}\to\omega\pi^{0})=(1.16\pm 0.44_{\rm stat}\pm 0.16_{\rm syst})\times 10^{-4}$ by BESIII ar2407-20551 hinting possible contribution of $b_{1}(1235)^{+}\to\omega\pi^{+}$ for $D^{+}\to K^{0}_{S}\pi^{+}\omega$ decay. The branching fraction for the decay $D^{0}\to K^{-}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ is predicted to be $(2.86\pm 0.70)\times 10^{-3}$ in this work. Comparing with the measurement $(3.392\pm 0.044_{\rm stat}\pm 0.085_{\rm syst})\%$ for three-body decay $D^{0}\to K^{-}\pi^{+}\omega$ in Ref. prd105.032009 , one could conclude that the contribution of $\rho(770\&1450)^{+}\to\omega\pi^{+}$ is the dominated in the three-body decay $D^{0}\to K^{-}\pi^{+}\omega$.
@@ IV Summary
With the parameters $g_{\rho\omega\pi}=16.0\pm 2.0$ GeV-1 and $A_{1}=0.171\pm 0.036$ for the vector form factor $F_{\omega\pi}$, we predicted the branching fractions for the first time for the quasi-two-body decays $D_{s}^{+}\to\eta[\rho^{+}\to]\omega\pi^{+}$, $D^{+}\to K^{0}_{S}[\rho^{+}\to]\omega\pi^{+}$ and $D^{0}\to K^{-}[\rho^{+}\to]\omega\pi^{+}$ with the intermediate $\rho^{+}=\{\rho(770)^{+},\rho(1450)^{+},\rho(770)^{+}+\rho(1450)^{+}\}$. By comparing our predictions with the experimental data, we found that the contributions from the subprocess $\rho(770)^{+}\to\omega\pi^{+}$ are significant in these three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$, notwithstanding the contributions originating from the Breit-Wigner tail effect of $\rho(770)^{+}$. The interference effects between the resonances $\rho(770)^{+}$ and $\rho(1450)^{+}$ make the $P$-wave resonance contributions for $\omega\pi^{+}$ pair enhanced in these three-body decays. Our numerical results suggest that the Cabibbo-favored three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$ are dominated by the $P$-wave intermediate states $\rho(770)^{+}$, $\rho(1450)^{+}$ and their interference effects.

### v2

@@ III Numerical results and Discussions
The key input in this work is the coupling constant $g_{\rho\omega\pi}$ in Eq. (21) for the subprocess $\rho(770)\to\omega\pi$; its value can be estimated with the relation $g_{\rho\omega\pi}\approx 3g_{\rho\pi\pi}^{2}/(8\pi^{2}F_{\pi})$ prc83.048201 with $F_{\pi}=92$ MeV PDG-2024 and can also be calculated from the decay width of $\omega\to\pi^{0}\gamma$ PDG-2024 . In the numerical calculation of this work, we adopt the value $g_{\rho\omega\pi}=16.0\pm 2.0$ GeV-1 JHEP2401-047 by taking into account the corresponding fitted and theoretical values in Refs. prd94.112001 ; prd86.057301 ; prd61.072003 ; JETPL94-734 ; prd55.249 ; prd77.113011 ; prd86.037302 ; prd88.054013 for it. In view of the expression of $F_{\omega\pi}(s)$, we have a constraint JHEP2401-047
@@
Table 3: The predicted branching fractions for the relevant quasi-two-body $D_{(s)}$ decays with the subprocesses $\rho(770,1450)^{+}\to\omega\pi^{+}$ along with the experimental measurements in PDG-2024 for the related three-body decay processes.
@@
    Three-body modes   Data PDG-2024
  $D_{s}^{+}\to\eta\omega\pi^{+}$   $(5.4\pm 1.3)\times 10^{-3}$
  $D^{+}\to K^{0}_{S}\omega\pi^{+}$   $(7.1\pm 0.5)\times 10^{-3}$
  $D^{0}\to K^{-}\omega\pi^{+}$   $(3.39\pm 0.10)\times 10^{-2}$
@@
The absolute branching fraction for the three-body decay $D^{+}\to K^{0}_{S}\pi^{+}\omega$ was determined to be $(0.707\pm 0.041_{\rm stat}\pm 0.029_{\rm syst})\%$ by BESIII in Ref. prd105.032009 , which is $4.6$ times larger than the prediction $(1.53\pm 0.37)\times 10^{-3}$ in Table 3 for the quasi-two-body decay $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$. We need to stress that the contribution weight of $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ in the corresponding three-body decay could be enhanced by improving the calculation method employed in this work, since our prediction for $D^{+}\to K^{0}_{S}[\rho(770)^{+}\to]\pi^{+}\pi^{0}$ in Table 2 is just around half of the data in PDG-2024 . Which also means the branching fraction for $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ could be twice of the corresponding prediction in Table 3. In addition, the $D^{+}\to b_{1}(1235)$ transition with a emitted $K^{0}_{S}$, could generate the $\omega\pi^{+}$ pair from $b_{1}(1235)$ for $D^{+}\to K^{0}_{S}\pi^{+}\omega$ decay. The $\omega\pi^{+}$ pair generated from the $D^{+}\to b_{1}(1235)$ transition is not like it from the matrix element $\langle\omega\pi|V^{\mu}\!-\!A^{\mu}|0\rangle$ in $D_{s}^{+}\to\eta\omega\pi^{+}$ decay, whose amplitude is just proportional to the mass difference between $u$ and $d$ quarks. The measurement of ${\mathcal{B}}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e})\times{\mathcal{B}}(b_{1}(1235)^{0}\to\omega\pi^{0})=(1.16\pm 0.44_{\rm stat}\pm 0.16_{\rm syst})\times 10^{-4}$ by BESIII ar2407-20551 hinting possible contribution of $b_{1}(1235)^{+}\to\omega\pi^{+}$ for $D^{+}\to K^{0}_{S}\pi^{+}\omega$ decay. But we know the contribution shouldn’t be large since we have the theoretical estimation $1.7\times 10^{-3}$ as the branching fraction for $D^{+}\to\bar{K}^{0}b_{1}(1235)^{+}$ prd67.094007 .
The input parameters for the quasi-two-body $D$ meson decays were updated in Ref. ar2503.18593 very recently. With those updated inputs in ar2503.18593 , the quasi-two-body decay branching fractions for the $D_{s}^{+}\to\eta\rho(770)^{+}$ and $D^{0}\to K^{-}\rho(770)^{+}$ agree well with these corresponding results in Tables 2-3 with the subprocesses $\rho(770)^{+}\to\pi^{+}\pi^{0}$ and $\rho(770)^{+}\to\omega\pi^{+}$, respectively. While the branching fractions for the quasi-two-body decays $D^{+}\to K^{0}_{S}[\rho(770)^{+}\to]\pi^{+}\pi^{0}$ and $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ will be increased to $(4.45\pm 0.14)\%$ and $(2.63\pm 0.64)\times 10^{-3}$, respectively. And the latter numerical value is about $37\%$ of the total branching fraction of $D^{+}\to K^{0}_{S}\pi^{+}\omega$ in prd105.032009 . For this three-body decay $D^{+}\to K^{0}_{S}\pi^{+}\omega$, one also has the contribution from $D^{+}\to\bar{K}^{*0}\pi^{+}$ with the subprocess $\bar{K}^{*0}\to K^{0}_{S}\omega$. When we take into account the branching fraction $(1.04\pm 0.12)\%$ for $D^{+}\to\bar{K}^{*0}\pi^{+}$ along with $\bar{K}^{*0}\to K^{-}\pi^{+}$ PDG-2024 and the discussions for the virtual contributions of $K^{*}$ in 2508.09578 , we can’t expect a large contribution from $D^{+}\to\bar{K}^{*0}\pi^{+}\to K^{0}_{S}\omega\pi^{+}$ for this three-body decay process. In addition, the cascade decay $D^{+}\to K^{*}(892)^{+}\omega\to K^{0}_{S}\pi^{+}\omega$ is a doubly Cabibbo suppressed model. Then we are pretty sure that the quasi-two-body process $D^{+}\to K^{0}_{S}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ should be the most important one for the three-body decay $D^{+}\to K^{0}_{S}\pi^{+}\omega$.
The branching fraction for the decay $D^{0}\to K^{-}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ is predicted to be $(2.86\pm 0.70)\times 10^{-3}$ in this work, which is less than a tenth of the measurement $(3.392\pm 0.044_{\rm stat}\pm 0.085_{\rm syst})\%$ for three-body decay $D^{0}\to K^{-}\pi^{+}\omega$ in Ref. prd105.032009 . The data $(6.5\pm 3.0)\times 10^{-3}$ for $D^{0}\to\bar{K}^{*}(892)^{0}\omega$ with the subprocesses $\bar{K}^{*}(892)^{0}\to K^{-}\pi^{+}$ and $\omega\to\pi^{+}\pi^{-}\pi^{0}$ is about twice of our result for $D^{0}\to K^{-}[\rho(770\&1450)^{+}\to]\omega\pi^{+}$ in this work. Actually, the three-body decay $D^{0}\to K^{-}\pi^{+}\omega$ is very different from the process $D_{s}^{+}\to\eta\omega\pi^{+}$, the former has very rich intermediate states. One has the resonances $\bar{K}^{*}(892,1410,1680)^{0},\bar{K}_{0,2}^{*}(1430)^{0}$, ${K}^{*}(892,1410,1680)^{-},{K}_{1}(1270)^{-}$ and $\rho(770,1450)^{+}$ decaying into $K^{-}\pi^{+}$, $K^{-}\omega$ and $\pi^{+}\omega$, respectively, in this three-body decay $D^{0}\to K^{-}\pi^{+}\omega$. The analyses of the complete resonance contributions for the decay process $D^{0}\to K^{-}\pi^{+}\omega$ is beyond the scope of this work, we shall leave them to the future studies.
@@ IV Summary
With the parameters $g_{\rho\omega\pi}=16.0\pm 2.0$ GeV-1 and $A_{1}=0.171\pm 0.036$ for the vector form factor $F_{\omega\pi}$, we predicted the branching fractions for the first time for the quasi-two-body decays $D_{s}^{+}\to\eta[\rho^{+}\to]\omega\pi^{+}$, $D^{+}\to K^{0}_{S}[\rho^{+}\to]\omega\pi^{+}$ and $D^{0}\to K^{-}[\rho^{+}\to]\omega\pi^{+}$ with the intermediate $\rho^{+}=\{\rho(770)^{+},\rho(1450)^{+},\rho(770)^{+}+\rho(1450)^{+}\}$. By comparing our predictions with the experimental data, we found that the contributions from the subprocess $\rho(770)^{+}\to\omega\pi^{+}$ are significant in these three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$, $D^{+}\to K^{0}_{S}\omega\pi^{+}$ and $D^{0}\to K^{-}\omega\pi^{+}$, notwithstanding the contributions originating from the Breit-Wigner tail effect of $\rho(770)^{+}$. The interference effects between the resonances $\rho(770)^{+}$ and $\rho(1450)^{+}$ make the $P$-wave resonance contributions for $\omega\pi^{+}$ pair enhanced in these three-body decays. The numerical results of this study suggest that the dominant resonance contributions for the three-body decays $D_{s}^{+}\to\eta\omega\pi^{+}$ and $D^{+}\to K^{0}_{S}\omega\pi^{+}$ are originated from the $P$-wave intermediate states $\rho(770)^{+}$, $\rho(1450)^{+}$ decaying into $\omega\pi^{+}$ pair and their interference effects.